20. Ramkali saved Rs 5 in the first week of a year and then increased her weekly saving by Rs 1.75. If in the nth week, her weekly savings become Rs 20.75, find n.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Given that, Ramkali saved Rs.5 in first week and then started saving each week by Rs.1.75. Hence, First term, a = 5 and common difference, d = 1.75 Also given, an = 20.75 Find, n = ? As we know, by the nth term formula, an = a+(n−1)d Therefore, 20.75 = 5+(n -1)×1.75 15.75 = (n -1)×1.75 (n -1) = 15.75/1.75 = 1575/175 = 63/7 = 9 n -1 = 9 n = 10 Hence, n is 10.